How many subsets of the set {1, 2, 3, 4, 5, 6, 7, 8} do not contain two consecutive integers ?


1
Expert's answer
2021-10-17T16:39:26-0400

Let S = {1, 2, . . . , 8} and let "S_e=\{2,4,6,8\}" and "S_o=\{1,3,5,7\}" .

A subset T of S has no element which are consecutive integers if and only if "T=T_o\subset T_e" where "T_e \subset S_e"

and "T_o\subset S_o" , such that Te has no two consecutive elements from the list 2, 4, . . . ,8 and To

contains no two consecutive elements of the list 1, 3, . . . , 7. Furthermore if an allowable

subset T of S is given, then the sets Te and To are uniquely determined, and Te and To

satisfy the stated conditions. Thus if "a_4" is the number of ways to form a set from the

list 2, 4, 6, 8 containing no two consecutive elements, then the number of subsets of

S with no elements differing by 2 is "a_4^2" We determine "a_4" .

To simplify notation, let "a_n" be the number of subsets of Sn = {1, 2, . . . , n} containing no

two consecutive element of Sn. It is easy to see that "a_1" = 2 (we get the empty set φ and

the set {1}) and "a_2=3" . Now suppose that n ≥ 3. If an allowable subset of Sn contains

n, then the other elements of the set can make up any allowable subset of "S_{n-2}" . If the

allowable subset does not contain n, then the elements of the set must be an allowable

subset of "S_{n-1}" . Thus

"a_n=F_{n+2}" where "F_n" denotes the "n^{th}" Fibonacci number. (the Fibonacci numbers are


defined by "F_0=0,F_1=1" and "F_n=F_{n-1}+F_{n-2}" for n ≥ 2.) Thus we find that the number

of subsets of "S_8" with no element which are consecutive integer

"a_4^2=F_{4+2}^2=F_6^2=8^2=64"